OpenStudy (anonymous):

Subtract 2/4+sqrt6 - 2/4-sqrt6

OpenStudy (anonymous):

\[\frac{ 2 }{ 4+\sqrt{6} } - \frac{ 2 }{ 4 -\sqrt{6} }\]